# Understanding Language Prior of LVLMs by Contrasting Chain-of-Embedding

### v2

@@ 5 Theoretical Analysis
We provide a new interpretation for our metric, $\mathbf{D}_{l}(\mathcal{P}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{P}_{\text{T}},F_{\theta})$, that locates VIP (Theorem 5.1) and how we can practically employ the expected representation distance (Theorem 5.2) through theoretical analyses. All the proofs and an additional theorem that justifies the use of our empirical representation distance (Lemma F.1) are given in Appendix F.
@@
By recasting the representation distance measurement as a density estimation problem, i.e., $d(Z^{l}_{\text{vis}},Z^{l}_{\text{blind}})\propto-\log\hat{p}_{\text{T}}(Z^{l})$ (please see Lemma F.2), we show that the difference in expected representation distances, $\mathbf{D}_{l}(\mathcal{P}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{P}_{\text{T}},F_{\theta})$, which we call representation divergence here, can be interpreted as a relative distributional discrepancy that measures how far our density estimator $\hat{p}_{\text{T}}(Z^{l})$, defined by $d(Z^{l}_{\text{vis}},Z^{l}_{\text{blind}})$, from a population distribution $p_{\text{VT}}(Z^{l})$ compared to $p_{\text{T}}(Z^{l})$ in Theorem 5.1.
@@
Implication. Lemma F.2 shows that $d(Z_{\text{vis}},Z_{\text{blind}})$ becomes a density estimator on $\mathcal{P}_{\text{T}}$ with a strictly proper scoring. With that, Theorem 5.1 tells us that $\mathbf{D}_{l}(\mathcal{P}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{P}_{\text{T}},F_{\theta})$ can be interpreted as a relative proximity of our density estimate $\hat{p}_{\text{T}}$ to each distributions $p_{\text{VT}}$ and $p_{\text{T}}$ with an additive constant $\bar{\mathbf{H}}$, defined by $p_{\text{VT}}$ and $p_{\text{T}}$. More intuitively, the first term, $\text{KL}(p_{\text{VT}}||\hat{p}_{\text{T}})$, can be understood how the ground truth representation distribution on $\mathcal{P}_{\text{VT}}$ far from $\hat{p}_{\text{T}}$ (distance-based estimate of $\mathcal{P}_{\text{T}}$), while the second term, $\text{KL}(p_{\text{T}}||\hat{p}_{\text{T}})$, can be regarded as a quality of our density estimation with $\hat{p}_{\text{T}}$ to approximate the distribution $p_{\text{T}}$.
@@
Implication. The first inequality (Ineq. 6) reveals a relationship between two expected representation distances across $\mathcal{P}_{\text{VT}}$ and $\mathcal{D}_{\text{T}}\sim\mathcal{P}_{\text{T}}$ with $d_{\bar{\mathcal{H}}}(\mathcal{D}_{\text{VT}},\mathcal{D}_{\text{T}})$ as a bridge. This tells us that if we want to increase visual integration after VIP for an unknown data distribution that require visual reasoning ($\mathcal{P}_{\text{VT}}$), we can pursue a greater lower bound of it by decreasing $\mathbf{D}_{l}(\mathcal{D}_{\text{T}},F_{\theta})$ and $d_{\bar{H}}(\mathcal{D}_{\text{VT}},\mathcal{D}_{\text{T}})$ with empirical samples we have. Meanwhile, in a case where we encountered an unknown mixture distribution $\mathcal{P}_{\text{M}}$, the second inequality (Ineq. 7) tells us we can broaden the effective range of $\mathbf{D}_{l}$ on $\mathcal{P}_{\text{M}}$ by pursuing greater value of $d_{\bar{\mathcal{H}}}(\mathcal{D}_{\text{VT}},\mathcal{D}_{\text{T}})$.
@@ 6 Conclusion
In this work, we present a formal framework for understanding and quantifying the language prior in LVLMs by contrasting the chain-of-embedding between visual and blind contexts. Through this framework, we identify the consistent existence of the Visual Integration Point (VIP), a specific layer at which the model begins to meaningfully incorporate visual context for task-solving beyond the shallow information gathering. Building on this observation, we propose a new metric named Total Visual Integration (TVI), which estimates the degree of effective visual integration and therefore language prior. We conduct comprehensive experiments across 9 LVLMs and 6 datasets, and the results demonstrate that our framework robustly works across models and tasks, providing consistent and interpretable signals about the presence and strength of language prior. Finally, we present some theorems for better understanding and justification of our framework. We hope that this work sheds light on the internal mechanisms of multimodal models and provides a foundation for diagnosing the language prior, ultimately guiding the development of reliable and responsible LVLMs.
@@ Acknowledgement
The authors would like to thank Boyi Li, Haobo Wang, Junbo Zhao, Min-Hsuan Yeh, and Jiatong Li for their insightful feedback and helpful discussions throughout the development of this work. Their suggestions greatly contributed to improving the quality of the draft. Changdae Oh, Seongheon Park, and Sharon Li are supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon.
Ethics Statement
This work conducts empirical and analytical studies on the internal behavior of LVLMs, with the goal of understanding and quantifying their reliance on language priors and the extent of visual information integration during inference. To pursue high standards of scientific excellence, we propose a formal framework with clear definitions of all used terms and conduct validation at scale, e.g., 54 combinations of models and datasets, and we further provide theoretical analyses on our framework. Our study does not involve any human subjects, personally identifiable information, or sensitive data. All experiments are conducted using publicly available models and benchmark datasets that are widely adopted in the multimodal learning community. Our proposed metrics and analyses are intended for research and diagnostic purposes. By providing tools to diagnose when LVLMs rely on text versus vision, we aim to support more accountable model development and contribute positively to the responsible advancement of AI. We encourage future work to further validate these findings under more diverse real-world conditions.
Reproducibility Statement
@@
Figure 6: Complete experimental results. (Part 1)
Figure 7: Complete experimental results. (Part 2)
Figure 8: Complete experimental results. (Part 3)
@@ Appendix B Discussion on Visual Integration Point
B.1 Configuration and Practice
In Eq. 8, we chracterized the VIP $l^{*}$ based on the pre-$l^{*}$ and post-$l^{*}$ representation divergences $\mathbf{D}_{l}(\mathcal{P}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{P}_{\text{T}},F_{\theta})$ where the pre-$l^{*}$ chain-of-embeddings shows nearly zero representation divergence whereas the post-$l^{*}$ chain-of-embeddings exhibits an effectively large representation divergence defined by positive $\tau$. In practice, we can not access the population distribution $\mathcal{P}_{\text{VT}}$ and $\mathcal{P}_{\text{T}}$, therefore we can not compute the truth expected representation distance $\mathbf{D}_{l}(\mathcal{P}_{\star},F_{\theta})$. What we do in practice is estimate that expected representation distance with finite samples $\mathcal{D}_{\text{VT}}$ and $\mathcal{D}_{\text{T}}$. We observed that this empirical estimator (Eq. 2) is actually well aligned with our hypothesis, yet there can be some cases where the fitness of the estimator is bad, e.g., Figure 9. However, even in that case, the point $l^{*}$, where the empirical representation divergence exceeds a positive constant for all subsequent layers, consistently emerges, and the TVI (Eq. 4) is calculated based on that $l^{*}$ becomes a strong indicator of LP.
@@ B.2 Algorithmic Estimation on Visual Integration Point
As discussed in §3.2, identifying the VIP provides valuable insights into when visual information begins to influence answer decoding. In most cases, we rely on empirical observation of the representation distance curves to determine $l^{*}$, which proves to be straightforward and interpretable across a wide range of models. However, in situations where manual inspection is impractical — e.g., for large-scale model comparisons or automated pipelines — it may be desirable to estimate the VIP programmatically. To this end, we introduce an optional empirical estimator that can automatically approximate the VIP given a model and a dataset. While not required for our core analysis, this estimation method serves as a practical tool in settings where manual identification of $l^{*}$ is infeasible.
Definition B.1 (Visual integration point estimator).
For a dataset $\mathcal{D}$ sampled from a mixture distribution $\mathcal{P}=(1-\lambda)\mathcal{P}_{\text{VT}}+\lambda\mathcal{P}_{\text{T}}$ for $\lambda\in(0,1)$ and an LVLM $F_{\theta}$ with $L$ decoder layers which produces a chain-of-embedding $(Z^{1},...,Z^{L})$ given input, let $\mathcal{L}$ be the index set of LVLM decoder layers and $\mathbf{D}_{l}$ be an expected representation distance defined as Eq. 2. An empirical visual integration point estimator is as follows,
$\text{VIP}(\mathcal{D},F_{\theta})$ $=\arg\min_{l\in\mathcal{L}\backslash L}\big[\mathbf{D}_{l}(\mathcal{D}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{D}_{\text{T}},F_{\theta})\geq\alpha\cdot\frac{\sigma_{l}}{\mathbf{D}_{l}(\mathcal{D},f)}\big],$ (8)
where $\alpha$ is a pre-defined coefficient and $\sigma_{l}$ denotes the sample standard deviation of the observed representation distance over $\mathcal{D}$.
$\text{VIP}(\mathcal{D},F_{\theta})$ can be interpreted as the layer index where the average $d(\cdot,\cdot)$ on $\mathcal{D}_{\text{VT}}$ first exceeds that on $\mathcal{D}_{\text{T}}$ up to an empirical coefficient of variation (Everitt & Skrondal, 2010) of $d(\cdot,\cdot)$ on $l$-th layer. Besides, one can also devise another VIP estimator based on the test statistic we discuss in Eq 10 of Lemma F.1 by specifying arbitrary significance levels that the user prefers.
@@ Appendix C Implementation Details
All TVI values reported in our experiments are computed based on empirically determined Visual Integration Points (VIPs) specific to each model. The following VIPs are used: Qwen2.5-VL-7B (18), InternVL3-8B (16), Gemma-3-4B (20), Gemma-3-12B (26), Gemma-3-27B (35), LLaVA-v1.5-7B (9), Eagle2.5-8B (15), Llama-3.2-11B-Vision (12), LLaVA-NeXT-Vicuna-7B (12), LLaVA-OV-Qwen2-7B (15) and SmolVLM (15). We also provide an automatic method for estimating VIP in Appendix B.2 for potential practical usage. Representation distances are computed using the hidden states corresponding to the last generated token. The metrics introduced in §4 are computed as follows:
@@
Appendix D Additional Experimental Results
@@ Appendix E Further Analysis
In this section, we elaborate on the limitations of attention-based and output-based analysis for language prior in LVLMs.

### v3

@@ 5 Theoretical Analysis
Next, we introduce a new interpretation for our measure, $\mathbf{D}_{l}(\mathcal{P}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{P}_{\text{T}},F_{\theta})$, that locates VIP (Theorem 5.1) and discuss how we can practically employ the expected representation distance (Theorem 5.2) through theoretical analyses. All the proofs and an additional theorem that justifies the use of our empirical representation distance (Lemma F.1) are given in Appendix F.
@@
By recasting the representation distance measurement as a density estimation problem, i.e., $d(Z^{l}_{\text{vis}},Z^{l}_{\text{blind}})\propto-\log\hat{p}_{\text{T}}(Z^{l})$ (please see Lemma F.2), we show that the difference in expected representation distances, $\mathbf{D}_{l}(\mathcal{P}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{P}_{\text{T}},F_{\theta})$, which we call representation divergence here, can be interpreted as a relative distributional discrepancy that measures how far the density estimator $\hat{p}_{\text{T}}(Z^{l})$, defined by $d(Z^{l}_{\text{vis}},Z^{l}_{\text{blind}})$, from a population distribution $p_{\text{VT}}(Z^{l})$ compared to $p_{\text{T}}(Z^{l})$ in Theorem 5.1.
@@
Implication. Theorem 5.1 tells us that $\mathbf{D}_{l}(\mathcal{P}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{P}_{\text{T}},F_{\theta})$ can be interpreted as a relative proximity of the density estimate $\hat{p}_{\text{T}}$ to each distributions $p_{\text{VT}}$ and $p_{\text{T}}$ with an additive constant $\bar{\mathbf{H}}$. Intuitively, the first term, $\text{KL}(p_{\text{VT}}||\hat{p}_{\text{T}})$, can be understood how $\hat{p}_{\text{T}}$ (estimate of $p_{\text{T}}$) far from the true representation distribution on $\mathcal{P}_{\text{VT}}$ while the second term, $\text{KL}(p_{\text{T}}||\hat{p}_{\text{T}})$, is a quality of the density estimation with $\hat{p}_{\text{T}}$ to approximate $p_{\text{T}}$. This expression converts the expected representational distance of the LVLM $F_{\theta}$ over $p_{\text{VT}}$ and $p_{\text{T}}$ into an information-theoretic divergence, the amount of surprise if we approximate the distribution over $p_{\text{VT}}$ via a blind-representation-centered Gaussian estimator $\hat{p}_{\text{T}}(Z^{l})=\mathcal{N}(\cdot;Z^{l}_{\text{blind}},I)$.
@@
Implication. The first inequality (Ineq. 7) reveals a relationship between two expected representation distances across $\mathcal{P}_{\text{VT}}$ and $\mathcal{D}_{\text{T}}\sim\mathcal{P}_{\text{T}}$ with $d_{\bar{\mathcal{H}}}(\mathcal{D}_{\text{VT}},\mathcal{D}_{\text{T}})$ as a bridge. This tells us that if we want to increase visual integration for an unknown data distribution that require visual reasoning ($\mathcal{P}_{\text{VT}}$), we can pursue a greater lower bound of it by decreasing $\mathbf{D}_{l}(\mathcal{D}_{\text{T}},F_{\theta})$ and $d_{\bar{H}}(\mathcal{D}_{\text{VT}},\mathcal{D}_{\text{T}})$ with empirical samples we have. Meanwhile, in a case where we encountered an unknown mixture distribution $\mathcal{P}_{\text{M}}$, the second inequality (Ineq. 8) tells us we can broaden the effective range of $\mathbf{D}_{l}$ on $\mathcal{P}_{\text{M}}$ by pursuing greater value of $d_{\bar{\mathcal{H}}}(\mathcal{D}_{\text{VT}},\mathcal{D}_{\text{T}})$.
@@ 6 Conclusion
In this work, we present a formal framework for understanding and quantifying the language prior in LVLMs by contrasting the chain-of-embedding between visual and blind contexts. Through this framework, we identify the consistent existence of the Visual Integration Point (VIP), a specific layer at which the model begins to meaningfully incorporate visual context for task-solving beyond the shallow information gathering. Building on this observation, we propose a new metric named Total Visual Integration (TVI), which estimates the degree of effective visual integration and therefore language prior. We conduct comprehensive experiments across 9 LVLMs and 6 datasets, and the results demonstrate that our framework robustly works across models and tasks, providing consistent and interpretable signals about the presence and strength of language prior. Finally, we present some theorems for better understanding and utilization of our framework. We hope that this work sheds light on the internal mechanisms of multimodal models and provides a foundation for diagnosing the language prior, ultimately guiding the development of reliable and responsible LVLMs.
Limitations.
We set the language prior to LVLMs as our sole target of analysis here, and developed our method based on the representation dynamics of LVLMs. However, there are many other potential biases and vulnerabilities originating from query distribution shifts in the wild, which may induce remarkable changes in the representation space and thus degradation of downstream performance (Verma et al., 2024; Oh et al., 2025a; b; Kim et al., 2025). Reliability of TVI-based language prior estimation should be further validated under realistic distribution shifts.
Besides, our method requires white-box access to the model’s internal hidden states and attention patterns. This restricts its applicability to open-weight models and excludes commercial APIs or closed-source systems. However, our framework is primarily designed for model analysis and interpretability research of white-box models, rather than serving as a versatile tool, aiming to shed light on how and when visual information is integrated during inference.
@@ Acknowledgement
The authors would like to thank Boyi Li, Haobo Wang, Junbo Zhao, Samuel Yeh, and Jiatong Li for their insightful feedback and helpful discussions throughout the development of this work. Their suggestions greatly contributed to improving the quality of the draft. Changdae Oh, Seongheon Park, and Sharon Li are supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon.
@@
Figure 6: Complete experimental results. (Part 1) LLaVA-v1.5-7B, LLaVA-NeXT-Vicuna-7B, and LLaVA-OV-Qwen2-7B on six datasets.
Figure 7: Complete experimental results. (Part 2) Gemma-3-4B, SmolVLM, and Llama-3.2-11B-Vision on six datasets.
Figure 8: Complete experimental results. (Part 3) Qwen2.5-VL-7B, Eagle2.5-8B, and InternVL3-8B on six datasets.
@@ Appendix B Discussion on Visual Integration Point
B.1 Note on VIP determination
In Eq. 3, we defined the VIP $l^{*}$ based on the pre-$l^{*}$ and post-$l^{*}$ representation divergences $\mathbf{D}_{l}(\mathcal{P}_{\text{VT}},F_{\theta})-\mathbf{D}_{l}(\mathcal{P}_{\text{T}},F_{\theta})$ where the pre-$l^{*}$ chain-of-embeddings shows nearly zero representation divergence whereas the post-$l^{*}$ chain-of-embeddings exhibits an effectively large representation divergence defined by positive $\tau$. In practice, we can not access the population distribution $\mathcal{P}_{\text{VT}}$ and $\mathcal{P}_{\text{T}}$, therefore we can not compute the truth expected representation distance $\mathbf{D}_{l}(\mathcal{P}_{\star},F_{\theta})$. What we do in practice is estimate that expected representation distance with finite samples $\mathcal{D}_{\text{VT}}$ and $\mathcal{D}_{\text{T}}$, and see the evolution of representation distance gap to manually pick the VIP $l^{*}$. We observed that this empirical estimator of representation distance (Eq. 2) works well as a measure for determining VIP in general, yet there can be some cases where the fitness of the estimator is bad, e.g., Figure 10. However, even in that case, the point $l^{*}$, where the empirical representation divergence exceeds a positive constant for all subsequent layers, consistently emerges, and the TVI (Eq. 11) is calculated based on that $l^{*}$ becomes a strong indicator of LP.
@@ B.2 Algorithmic Estimation on Visual Integration Point
As discussed in §3.2, identifying the VIP provides valuable insights into when visual information begins to influence answer decoding. In most cases, we rely on empirical observation of the representation distance curves to manually determine $l^{*}$, which proves to be straightforward and interpretable across a wide range of models. However, in situations where manual inspection is impractical — e.g., for large-scale model comparisons or automated pipelines — it may be desirable to estimate the VIP in a data-driven, automatic way. To this end, we introduce an algorithmic rule that can automatically estimate the VIP given a model and a dataset. While not required for our core analysis, this estimation method serves as a practical tool in settings where manual identification of $l^{*}$ is infeasible.
Below is an estimation strategy based on the test statistic we discussed in Eq 12 of Lemma F.1 by specifying arbitrary significance levels that a user prefers. To be specific, given a pooled sample standard deviation $\hat{\sigma}=\sqrt{\frac{\sigma^{2}_{l,{\text{T}}}}{|\mathcal{D}_{\text{T}}|}+\frac{\sigma^{2}_{l,{\text{VT}}}}{|\mathcal{D}_{\text{VT}}|}}$, we can define the estimated VIP as follows,
[…]
where $\sigma^{2}_{l,\text{T}}=\frac{\sum_{z\in\mathcal{D}_{T}}(d(z^{l}_{\text{vis}},z^{l}_{\text{blind}})-\mathbf{D}_{l}(\mathcal{D}_{\text{T}},F_{\theta}))^{2}}{|\mathcal{D}_{\text{T}}|-1}$ is a sample variance on $\mathcal{D}_{\text{T}}$ and $\sigma^{2}_{l,\text{VT}}$ is similarly defined. Intuitively, the quantity on the left-hand side denotes a deviation-normalized representation distance in the layer $l$, whereas the right-hand side means the historical average of those distances with a weighting coefficient $\beta$, which was set to 2.0 in our case. The above algorithm can be applied to any given dataset, and we conduct comprehensive experiments to show its robust performance across different models and datasets in Table 6 and Figure 9.
Table 6: Comparison of VIP detection methods and their effectiveness (Spearman correlations between the resulting TVI and downstream correctness).
Model (Layer #) VIP Detection MMBench - VIP ($\rho$) ViLP - VIP ($\rho$)
Qwen2.5-VL-7B (28) Manual 18 (0.6335) 18 (0.6335)
Qwen2.5-VL-7B (28) Variance-based 18 (0.6335) 19 (0.6336)
Gemma-3-4B (30) Manual 20 (0.7970) 20 (0.7970)
Gemma-3-4B (30) Variance-based 16 (0.7973) 16 (0.7973)
InternVL3-8B (32) Manual 16 (0.5709) 16 (0.5709)
InternVL3-8B (32) Variance-based 17 (0.5749) 20 (0.5949)
Figure 9: Comparison between the manually selected VIPs and automatically selected VIPs.
From Table 6 and Figure 9, we can see that VIP estimated by the proposed algorithm are quite close to the result of the manually selected one and clearly mark where the two curves start to diverge (based on the divergence plots), and the Spearman correlations between TVIs computed by those estimated VIPs and the downstream prediction correctness are robust across different datasets.
@@ Appendix C Implementation Details
All TVI values reported in our experiments are computed based on empirically determined Visual Integration Points (VIPs) specific to each model. The following VIPs are used: Qwen2.5-VL-7B (18), InternVL3-8B (16), Gemma-3-4B (20), Gemma-3-12B (26), Gemma-3-27B (35), LLaVA-v1.5-7B (9), Eagle2.5-8B (15), Llama-3.2-11B-Vision (12), LLaVA-NeXT-Vicuna-7B (12), LLaVA-OV-Qwen2-7B (15) and SmolVLM (15) [^6] It should be noted that these manually selected VIPs are not necessarily optimal; however, they already achieve strong and robust effectiveness and are sufficient for analytical purpose. The truly optimal VIP is likely to lie in their vicinity.. We also provide an automatic method for estimating VIP in Appendix B.2 for potential practical usage. Representation distances are computed using the hidden states corresponding to the last generated token. The metrics introduced in §4 are computed as follows:
@@
Appendix D Full Results of the Main Experiment
@@
Table 8: Summary of evaluated models and their architectural specifications.
Model Fusion # Layers Hidden Size
Qwen2.5-VL-7B/32B/72B MLP 28 / 64 / 80 3584 / 5120 / 8192
Gemma-3-4B/12B/27B Q-Former 34 / 48 / 62 2560 / 3840 / 5376
InternVL3-8B MLP 28 3584
LLaVA-v1.5-7B MLP 32 4096
Eagle2.5-8B MLP 28 3584
Llama-3.2-11B-Vision X-Attention 40 4096
LLaVA-NeXT-Vicuna-7B MLP 32 4096
LLaVA-OV-Qwen2-7B MLP 28 3584
SmolVLM MLP 24 2048
InstructBLIP-Vicuna-7B Q-Former 32 4096
Figure 11: Extended experiments across models of varying scales and fusion architectures.
Additional architectures & scales. We further validate our findings on models with diverse architectural designs, including Llama-3.2-11B-Vision with cross-attention–based multimodal fusion and InstructionBLIP-Vicuna-7B with a Q-Former–style fusion mechanism, as well as on larger-scale models such as Qwen2.5-VL-32B and Qwen2.5-VL-72B. As shown in Figure 11, the results consistently corroborate our conclusions across both architectural variants and model scales. Comprehensive statistics for all evaluated models are provided in Table 8.
@@ Appendix E Further Analysis
Analysis on the limitations of attention-based and output-based LP analysis.
@@
Ablation on aggregation strategies for TVI calculation.
In our main experiments, we adopt a simple aggregation strategy for computing TVI by averaging the representation distances across all post-VIP layers. To assess whether more sophisticated aggregation may improve the metric, we additionally experiment with a standard-deviation-based weighted TVI, defined as:
$\text{TVI}(l^{*};x,F_{\theta})=\frac{1}{L-l^{*}+1}\sum_{l=l^{*}}^{L}\big[\sigma_{l}\cdot d(z^{l}_{\text{vis}},z^{l}_{\text{blind}})\big]\text{,}$ (11)
Table 9: Comparison of aggregation methods for computing TVI.
Model Aggregation VLind ViLP
Qwen2.5-VL-7B Simple Average 0.7155 0.6335
Qwen2.5-VL-7B Std Reweighting 0.7164 0.6348
InternVL3-8B Simple Average 0.6727 0.5709
InternVL3-8B Std Reweighting 0.6739 0.5723
where $\sigma_{l}$ denotes the standard deviation of representation distances at layer $l$. This weighting scheme emphasizes layers whose distance distributions exhibit higher variability, i.e., layers that are more sample-specific and potentially more informative, while down-weighting contributions that may arise from less discriminative layers. As shown in Table 9, this reweighting provides a slight improvement in effectiveness over the simple averaging strategy across both VLind and ViLP. However, the gains are modest and come at the cost of additional computation. These findings suggest that our original simple averaging approach already serves as a strong and robust summary statistic, while the weighted variant may offer incremental refinement in specialized scenarios.
Analysis on instruction-level perturbation.
Figure 14: Impacts of instruction-level perturbation. (a) Representation distances obtained by contrasting chain-of-embedding sequences produced with and without textual instructions. (b) TVI scores under instructions with different styles
To strengthen our study of the contrasting chain-of-embedding, we compare embedding chains produced by image-text inputs versus image-only inputs (Figure 14 (a)). As expected, the two curves exhibit only negligible differences, likely due to the limited semantic content of image-only inputs. This phenomenon can be attributed to two main factors: (1) The hidden states from image-only inputs do not reveal anything about the model’s answer prediction, considering that the instructions are not even provided. The resulting representations thus are not directly comparable in the same behavioral space. (2) Most current LVLMs are not trained to handle image-only inputs for question answering, which makes the model’s behavior on image-only inputs unpredictable. Nonetheless, even under these constraints, we still observe that the revised TVI, which now specifically captures the contribution of textual content to the decoding process, is marginally higher for vision-independent samples than for vision-dependent ones. This further corroborates the robustness of our analysis framework and the consistency of our conclusions.
To further isolate the influence of the textual component in multimodal inputs, we investigate how different instruction styles affect model behavior. Specifically, we contrast misleading instructions (those originating from VLind-Bench or ViLP that elicit strong LP) with generic, vision-centric prompts such as “Describe the image in detail.” As results in Figure 14 (b) show, TVI under these vision-centric instructions is significantly higher than under misleading ones. This confirms that TVI is sensitive to the model’s bias toward language priors under different instruction regimes.
Analysis on VIP & TVI across different training stages.
To better understand how VIP and TVI evolve during model training, we analyze checkpoints from multiple stages of LLaVA-v1.5 visual instruction tuning. As shown in Figure 15 and Figure 16, our evaluation reveals two key observations. First, the VIP position remains remarkably stable across training stages (around the 12th layer). This suggests that the mechanism governing where visual information is integrated is largely established during pretraining and persists throughout subsequent finetuning. In other words, VIP appears to reflect an intrinsic architectural property rather than a behavior shaped by instruction tuning. Second, TVI exhibits a clear upward trend as training progresses, indicating that the degree to which the model effectively incorporates visual information improves gradually during instruction tuning. These findings align with our previous conclusion that VIP is a model-specific property as well as our expectation that visual instruction tuning incrementally enhances the model’s multimodal fusion capability.
Figure 15: Evaluation results across different stages of LLaVA-v1.5’s visual instruction tuning.
Figure 16: Average TVI on ViLP across different stages of visual instruction tuning.
Analysis on when TVI fails.
Althrough TVI shows strong correlation with answer correctness as shown in Table 3, there are still some non-neglectable portion on failure cases that TVI can not properly predict the model’s answer quality. Conceptually, TVI is designed to capture the degree of effective visual integration during decoding. While higher TVI generally implies that the generated answer is more grounded in visual input and thus more likely to be correct, there are natural regimes where this relationship weakens, for example: (1) Some questions are not strongly visually demanding; the model can answer them correctly with relatively little visual integration. In such cases, the answer may be correct even when TVI is small, leading to “false negatives” from TVI’s perspective. (2) For very challenging visual questions (e.g., fine-grained recognition, subtle spatial reasoning, or interpreting small/occluded objects) the model may still fail even after substantial effort to integrate visual information (high TVI). This yields “false positives” in terms of TVI–correctness alignment. These factors introduce inherent noise into any evaluation based on TVI vs. correctness.
To further characterize these patterns, we visualize the TVI distribution for ViLP samples that Qwen2.5-VL-7B answers correctly and perform a focused analysis on those with unexpectedly low TVI. As shown in Figure 17, while the majority of correctly answered samples exhibit relatively high TVI, a small subset attains low TVI despite being answered correctly. Upon inspection, many of these cases can indeed be solved without relying on visual input, either due to knowledge leakage or lucky guessing. This suggests that such questions are intrinsically less visually demanding, and the low TVI observed in these cases does not constitute a failure of the metric but instead reflects the fundamental properties of the task itself.